Calculate the change in enthalpy of the given reaction using Hess' Law

Given the following enthalpy changes:

H2 + 1/2 O2 -> H2O ΔH= -286kJ

CO2 -> C + O2 ΔH= +394 kJ

2 CO2 + H2O -> C2H2 + 5/2 O2 ΔH= +1300 kJ


Calculate the ΔH for the following reaction:

C2H2 -> 2C + H2


1
Expert's answer
2021-10-31T12:41:48-0400

H2O"\to" H2+1/2O2 "\Delta" H=+286kJ

2CO2"\to" 2C+2O2 "\Delta" H=2(394)kJ

C2H2 + 5/2 O2 "\to"2 CO2 + H2O  ΔH= -1300 kJ

ΔH=286+2(394)+(-1300)

=-226kJ